# SERA: Soft-Verified Efficient Repository Agents

### v1

@@ Abstract
Open-weight coding agents should hold a fundamental advantage over closed-source systems: they can be specialized to private codebases, encoding repository-specific information directly in their weights. Yet the cost and complexity of training has kept this advantage theoretical. We show it is now practical. We present Soft-Verified Efficient Repository Agents (SERA), an efficient method for training coding agents that enables the rapid and cheap creation of agents specialized to private codebases. Using only supervised finetuning (SFT), SERA achieves state-of-the-art results among fully open-source (open data, method, code) models while matching the performance of open-weight models like Devstral-Small-2. Creating SERA models is 26x cheaper than reinforcement learning and 57x cheaper than previous synthetic data methods to reach equivalent performance. Our method, Soft Verified Generation (SVG), generates thousands of trajectories from a single code repository. Combined with the cost-efficiency, this enables specialization to private codebases. SVG is built on two observations that emerged from simplification of previous methods: First, soft verification, where instead of testing the correctness of synthetic coding data via unit tests, we only compare the partial line-by-line overlap of patches generated from two rollouts. This removes the need for test infrastructure and enables data generation from any repository, practically removing limits on the amount of data we can generate from a single codebase as well as what codebases can be used. Second, vague instructions can diversify training data, increasing the proportion of data focused on non-bug related changes like refactoring. We find that these vague instructions improve SWE-bench performance as well as bug-focused data. In more detail, SVG is based on two rollouts from an agent: in the first, a teacher model is prompted with a vague instruction to make a change to a codebase starting from a randomly selected function, producing a trajectory and patch. This trajectory is converted into a synthetic pull request. In the second, the teacher model attempts to reproduce the patch given only the pull request description. Soft verification compares the two patches using line-level recall for training data selection. Taken together, this creates a cheap pipeline for high-quality data that enables rapid experimentation. We show through power scaling curves that private codebase specialization is highly sample efficient and matches or exceeds teacher model performance at low costs. Beyond repository specialization, we apply SVG to a larger corpus of codebases, generating over 200,000 synthetic trajectories. We use this dataset to provide detailed analysis of scaling laws, ablations, and confounding factors for training coding agents. Overall, we believe our work will greatly accelerate research on open coding agents and showcase the advantage of open-source models that can specialize to private codebases. We release SERA as the first model in Ai2’s Open Coding Agents series, along with all our code, data, and Claude Code integration to support the research community.
@@
[…]
@@ 1 Introduction
As the first release in Ai2’s Open Coding Agents series, our method trains a 32B coding agent with simple supervised finetuning achieving state-of-the-art open-source results at 40 GPU days (\$2000) or match strong open-weight models like Devstral-Small-2 at a budget of \$9000. When specializing to a particular codebase, our pipeline can match or exceed teacher model performance at \$1300.
@@
We found that much of the complexity in prior pipelines is unnecessary. Firstly, soft verification, where patches are checked by partial line-by-line matching rather than executed test suites, produces training data of equal quality to full test-based verification. At the scales we test, the degree of verification has minimal effect on downstream performance, removing the need for test infrastructure entirely and enabling data generation from any repository. This makes synthetic data generation much more straightforward. Secondly, some coding instructions are inherently vague, and we observe that models prompted to fix these issues often produce changes such as refactoring or documentation improvements that are more representative of real world tasks than just bug fixes. Rather than requiring bug-focused data, we find that this general coding data is equally effective to improve performance. Together, these findings mean that generating effective training data requires neither test infrastructure nor complex bug-injection pipelines. Additionally, unlike other methods, how much data we can generate from repositories are not limited by test coverage or quality.
@@ 2.1 SWE-bench
SWE-bench (Jimenez et al., 2023) is the standard benchmark for evaluating coding agents on real-world software engineering tasks. While other benchmarks exist that are more comprehensive, such as Terminal-Bench (Merrill et al., 2026), SWE-bench offers broader comparisons between methods and models, and its confounding factors are better understood. Each task is derived from a real GitHub issue and pull request from $12$ popular Python repositories such as Django, Sympy, and Sphinx. Given an issue description, the agent must produce a patch that resolves the issue. The repository’s test suite is run before and after applying the patch and a task is considered solved if previously failing tests now pass and no previously passing tests are broken. SWE-bench Verified is a curated subset where human annotators have verified that each task is solvable and that the tests correctly validate the solution. We use SWE-bench Verified for all evaluations in this work.
@@ 2.4 Reinforcement Learning
However, reinforcement learning introduces substantial infrastructure complexity. It requires online rollouts during training, sandboxed execution environments, and distributed systems for coordinating rollouts with gradient updates. This complexity is reflected in team sizes: recent RL papers for software engineering agents average 12 or more authors, including SWE-RL with 9 authors (Wei et al., 2025), SkyRL-Agent with 15 authors (Cao et al., 2025), and SWE-rebench with 9 authors (Badertdinov et al., 2025). Beyond team size, reinforcement learning is known to be highly unstable and difficult to use in practice, with training runs sensitive to hyperparameters, reward shaping, and random seeds (Henderson et al., 2018; Engstrom et al., 2020).
@@ 3.1 Soft Verified Generation (SVG)
Agent Workflow: We use SWE-agent (Yang et al., 2024) to generate trajectory rollouts. SWE-agent allows users to define a variety of tools available to an agent and gives users the ability to adjust settings such as the length of tool outputs and context history. To reduce the effect of confounding factors, we use SWE-agent in its vanilla state: we only provide the agent with the ability to run a view tool, edit tool, submission tool, and bash commands. Furthermore, we do not truncate context history or tool outputs at any point during rollouts. While truncations are frequently done to avoid context window errors, we noticed that many previous works use slightly different heuristics, making it difficult to objectively compare performance. Additionally, we believe an important trait of coding agents is their ability to solve tasks while avoiding unnecessarily long tool calls and outputs. However, for transparency, we ablate different heuristic choices in Appendix LABEL:.
First Rollout: At a high-level, the first rollout works as follows: We prompt the teacher model $\mathbf{M}$ with “There is a $\mathbf{bug_{j}}$ downstream of function $\mathbf{func_{i}}$.”, where $\mathbf{bug_{j}}$ is a high-level description of a bug type, $\mathbf{func_{i}}$ is a randomly chosen function in the codebase $\mathbf{C}$. The function $\mathbf{func_{i}}$ serves as an arbitrary starting point for the agent. We run the pipeline once for every function in the codebase. Each $\mathbf{bug_{j}}$ is randomly sampled from a larger list of 51 types of bugs $\mathbf{B}$ and asks the model to fix issues ranging from state management to code clarity. We generate this list of 51 bugs from papers that study bug distributions in software systems (Just et al., 2014; Widyasari et al., 2020). We intentionally leave the prompt vague to widen the range of acceptable changes. We rollout for a maximum of 115 steps, although this limit is rarely reached.
@@ 3.2 Training
We prioritize training on trajectories that are $\leq 32768$ tokens in length. To increase sample size as needed, we selectively truncate longer trajectories based on the ratio of trajectory steps within the context limit—we term this “truncation ratio”. In Section 5.2, we explore the effects of truncation in depth and explain why it must be done with caution.
@@ 4.1 Controlled Comparisons
We also note an important methodological consideration regarding context management during evaluation. Some agent frameworks employ optimizations such as retaining only the last few tool calls in context rather than the full trajectory history. While this compression allows models to appear effective at longer apparent context lengths, it introduces a confounding factor: methods using such optimizations may appear to benefit from increased context without the associated computational cost. Furthermore, such optimizations cause key-value cache invalidation during inference, which is prohibitively expensive for practical deployment. For fair comparison, we evaluate all methods using full context retention without such optimizations, ensuring that reported context lengths accurately reflect the actual information available to the model.
@@ 4.3 Repository Specialization
Initial Data Generation: Because every SWE-Bench instance is sourced from a unique commit, the set of instances from each repository will span multiple years. To account for this, we identify the earliest commit and latest commit in SWE-Bench Verified from each repository and generate data from five equally spaced commits in that period. While some functions are repeated across commits, each commit presents the codebase in a different context, which ensures trajectories in both the first and second rollout remain unique. Aggregating across commits, we obtain between $46000$ and $54000$ trajectories for each repository combined across both rollouts. We decide to train on both rollouts to increase sample size since the majority of the generated trajectories exceed $32768$ tokens. We investigate the effects of mixing rollouts in Section 5.5. Due to compute constraints, we train on 8000 trajectories per repository rather than the full dataset; however, we release all generated trajectories to enable future research to explore larger-scale specialization.
@@
Table 3: Specialization results at 32K context comparing GLM-4.5-Air (teacher) and fine-tuned Qwen 3-32B (student) on the three largest repositories in SWE-Bench Verified. Fine-tuned models are trained on 8000 synthetic trajectories from each repository. Results averaged over three seeds. Devstral-Small-2-24B results from Rastogi et al. (2025). See Table 13 in the appendix for 64K evaluation results.
@@
Specialization Scaling Law: To understand how data composition affects specialization efficiency, we fit scaling laws across different mixtures of Django-specific and general coding data (Figure 4). We define the specialization ratio $\alpha$ as the fraction of repository-specific data in the training mixture. At $\alpha=1.0$ (pure Django data), the model matches teacher performance (GLM-4.5-Air at 51.2%) with only 8,000 samples. In contrast, $\alpha=0.0$ (pure general data) is unable to reach equivalent performance even at 16,000 samples. Intermediate mixtures ($\alpha=0.75$, $\alpha=0.25$) show increasing asymptotic performance as the proportion of specialized data increases. This indicates that when training for a target codebase, the ratio of specialized data is the most important factor. Assumptions of ANOVA were met. A one-way ANOVA revealed a statistically significant effect of specialization ratio at 1,500 samples, $F(3,8)=10.78$, $p=.003$, $\eta^{2}=.80$. Post-hoc comparisons using Tukey’s HSD showed significant differences between $\alpha=1.0$ vs. $\alpha=0.0$ ($p=.009$) and $\alpha=0.75$ vs. $\alpha=0.0$ ($p=.005$). This indicates that repository-specific data yields better performance than general data at equivalent sample sizes.
@@ 5.1 Verification
In Figure 5, we ablate four different verification thresholds for $\mathbf{T_{2}}$ trajectories: $r=0.0$, $0.25$, $0.75$, and $1$. We also plot the performance of models trained on $\mathbf{T_{1}}$ trajectories (which are inherently unverified) at each scale for comparison. We study each verification threshold using only complete trajectories that fit within $32768$ tokens, with the final datapoint of each curve representing the largest possible dataset for that threshold. Only using trajectories $\leq 32768$ tokens reduces the total number of trainable trajectories but allows us to avoid confounding factors introduced by truncation.
If verification was essential for performance, we would expect to observe increasing performance as the verification threshold increases. Instead, scaling up to 7400 samples, we find that all verification thresholds perform similarly. For example, at the maximum scale, training on $\mathbf{T_{2}}$ trajectories that are soft-verified at $r=0.5$ shows no benefit over training on completely unverified $\mathbf{T_{2}}$ trajectories. Furthermore, models trained on $\mathbf{T_{1}}$ trajectories from the first rollout result in similar SWE-Bench Verified performance at each scale, despite representing a completely different distribution of coding tasks.
These results suggest that verification is not a necessity for high quality coding data, a behavior that is similar to what has been observed in other types of reasoning tasks (Chandra et al., 2026). We hypothesize that this is because even incorrect trajectories are can contain important skills, such as how to convert an intention into a relevant code edit even if the intention does not perfectly address the PR at hand. The assumption of normality was violated. A Kruskal-Wallis H-test revealed no statistically significant difference between verification thresholds, $H(3)=7.19$, $p=.066$, $\varepsilon^{2}=.52$. This confirms that soft verification performs as well as hard verification, and even unverified data achieves comparable results.
@@ 5.2 Truncation
Truncation is a popular but understudied practice across coding agent research to handle trajectories from teacher models whose context windows are larger than the base model’s. Without truncation, base model context limits make it impossible to train on significant amounts of data: $23.26\%$ of SWE-smith and $24.83\%$ of Sera-4.5A-Lite (verified at $r=1$) exceed $32768$ tokens.
@@
To test this hypothesis, we order $\mathbf{T_{1}}$ trajectories from Sera-4.5A-Lite based on the ratio of trajectory steps that fit in $32768$ tokens, a property we term “truncation ratio”. We partition the ordered $\mathbf{T_{1}}$ trajectories into subsets of 3000 samples each. This forces each subsequent partition to contain trajectories with strictly lower truncation ratios than the previous partition. We then train Qwen3-32B on each partition. $\mathbf{T_{1}}$ trajectories work well because they are longer than $\mathbf{T_{2}}$ trajectories on average while exhibiting similar scaling trends. This allows us to study the effect of training on a wide range of truncation ratios with a non-trivial amount of data and expect findings to translate.
In Figure 6 we plot SWE-Bench Verified performance against the average truncation ratio from every partition. Surprisingly, we find that the best data comes from trajectories that have high truncation ratios but are not fully contained in $32768$ tokens. Subsequent truncation ratios result in gradually decreasing performance. We suspect that this is due to a combination of factors, such as longer trajectories reflecting more difficult tasks and that a model’s final steps are typically focused on the redundant task of submitting its solution instead of problem solving.
@@
Assumptions were met. An independent samples t-test showed that ordered truncation ($M=43.00$, $SD=1.93$) significantly outperformed random truncation ($M=37.47$, $SD=0.50$), $t(4)=4.81$, $p=.009$, $d=3.93$. This confirms that the order in which content is truncated matters: preserving earlier turns in the trajectory is more effective than random truncation.

### v2

@@ Abstract
Open-weight coding agents should hold a fundamental advantage over closed-source systems: they can be specialized to private codebases, encoding repository-specific information directly in their weights. Yet the cost and complexity of training has kept this advantage theoretical. We show it is now practical. We present Soft-Verified Efficient Repository Agents (SERA), an efficient method for training coding agents that enables the rapid and cheap creation of agents specialized to private codebases. Using only supervised finetuning (SFT), SERA achieves state-of-the-art results among fully open-source (open data, method, code) models while matching the performance of open-weight models like Devstral-Small-2. Creating SERA models is 26x cheaper than reinforcement learning and 57x cheaper than previous synthetic data methods to reach equivalent performance. Our method, Soft Verified Generation (SVG), generates thousands of trajectories from a single code repository. Combined with cost-efficiency, this enables specialization to private codebases. SVG is built on two observations that emerged from simplification of previous methods: First, soft verification, where instead of testing the correctness of synthetic coding data via unit tests, we only compare the partial line-by-line overlap of patches generated from two rollouts. This removes the need for test infrastructure and enables data generation from any repository, practically removing limits on the amount of data we can generate from a single codebase as well as what codebases can be used. Second, vague instructions can diversify training data, increasing the proportion of data focused on non-bug related changes like refactoring. We find that these vague instructions improve SWE-bench performance as well as bug-focused data. In more detail, SVG is based on two rollouts from an agent: in the first, a teacher model is prompted with a vague instruction to make a change to a codebase starting from a randomly selected function, producing a trajectory and patch. This trajectory is converted into a synthetic pull request. In the second, the teacher model attempts to reproduce the patch given only the pull request description. Soft verification compares the two patches using line-level recall for training data selection. Taken together, this creates a cheap pipeline for high-quality data that enables rapid experimentation. We show through power scaling curves that private codebase specialization is highly sample efficient and matches or exceeds teacher model performance at low costs. Beyond repository specialization, we apply SVG to a larger corpus of codebases, generating over 200,000 synthetic trajectories. We use this dataset to provide detailed analysis of scaling laws, ablations, and confounding factors for training coding agents. Overall, we believe our work will greatly accelerate research on open coding agents and showcase the advantage of open-source models that can specialize to private codebases. We release SERA as the first model in Ai2’s Open Coding Agents series, along with all our code, data, and Claude Code integration to support the research community.
@@
Figure 1: (a) Scaling and cost comparison of coding agent training approaches using self-hosted vLLM inference. (b) Repository specialization scaling law on Django, where $\alpha$ denotes the fraction of Django-specific data in the training mixture. With full specialization ($\alpha=1.0$), the model matches teacher performance at 8k samples; general data alone ($\alpha=0.0$) is unable to match teacher performance, even with twice the sample size.
@@ 1 Introduction
As the first release in Ai2’s Open Coding Agents series, our method trains a 32B coding agent with simple supervised finetuning achieving state-of-the-art open-source results at 40 GPU days (\$2,000) or matching strong open-weight models like Devstral-Small-2 at a budget of \$9,000. When specializing to a particular codebase, our pipeline can match or exceed teacher model performance at \$1300.
@@
We found that much of the complexity in prior pipelines is unnecessary. Firstly, soft verification, where patches are checked by partial line-by-line matching rather than executed test suites, produces training data of equal quality to full test-based verification. At the scales we test, the degree of verification has minimal effect on downstream performance, removing the need for test infrastructure entirely and enabling data generation from any repository. This makes synthetic data generation much more straightforward. Secondly, some coding instructions are inherently vague, and we observe that models prompted to fix these issues often produce changes such as refactoring or documentation improvements that are more representative of real world tasks than just bug fixes. Rather than requiring bug-focused data, we find that this general coding data is equally effective to improve performance and can be produced by prompting a model on any repository in its initial state. Together, these findings mean that generating effective training data requires neither test infrastructure nor complex bug-injection pipelines. Additionally, unlike other methods, how much data we can generate from repositories is not limited by test coverage or quality.
@@ 2.1 SWE-bench
SWE-bench (Jimenez et al., 2023) is the standard benchmark for evaluating coding agents on real-world software engineering tasks. While other benchmarks exist that are more comprehensive, such as Terminal-Bench (Merrill et al., 2026), SWE-bench offers broader comparisons between methods and models, and its confounding factors are better understood. Each task is derived from a real GitHub issue and pull request from $12$ popular Python repositories such as Django, Sympy, and Sphinx. Given an issue description, the agent must produce a patch that resolves the issue. The repository’s test suite is run before and after applying the patch and a task is considered solved if previously failing tests now pass and no previously passing tests are broken. SWE-bench Verified is a curated subset where human annotators have verified that each task is solvable and that the tests correctly validate the solution. We use SWE-bench Verified for all evaluations.
@@ 2.4 Reinforcement Learning
However, reinforcement learning introduces substantial infrastructure complexity. It requires online rollouts during training, sandboxed execution environments, and distributed systems for coordinating rollouts with gradient updates. This complexity is reflected in team sizes: recent RL papers for software engineering agents average 12 or more authors, including SWE-RL with 9 authors (Wei et al., 2025), SkyRL-Agent with 15 authors (Cao et al., 2025), and SWE-rebench with 9 authors (Badertdinov et al., 2025). Beyond team size, reinforcement learning is known to be highly unstable and difficult to use in practice, with training sensitive to hyperparameters, reward shaping, and random seeds (Henderson et al., 2018; Engstrom et al., 2020).
@@ 3.1 Soft Verified Generation (SVG)
Agent Workflow: We use SWE-agent (Yang et al., 2024) to generate trajectory rollouts. SWE-agent allows users to define a variety of tools available to an agent and gives users the ability to adjust settings such as the length of tool outputs and context history. To reduce the effect of confounding factors, we use SWE-agent in its vanilla state: we only provide the agent with the ability to run a view tool, edit tool, submission tool, and bash commands. Furthermore, we do not truncate context history or tool outputs at any point during rollouts. While truncations are frequently done to avoid context window errors, we noticed that many previous works use slightly different heuristics, making it difficult to objectively compare performance. Additionally, we believe an important trait of coding agents is their ability to solve tasks while avoiding unnecessarily long tool calls and outputs.
First Rollout: At a high-level, the first rollout works as follows: We prompt the teacher model $\mathbf{M}$ with “There is a $\mathbf{bug_{j}}$ related to- function $\mathbf{func_{i}}$.”, where $\mathbf{bug_{j}}$ is a high-level description of a bug type, $\mathbf{func_{i}}$ is a randomly chosen function in the codebase $\mathbf{C}$. The function $\mathbf{func_{i}}$ serves as an arbitrary starting point for the agent. We run the pipeline once for every function in the codebase. Each $\mathbf{bug_{j}}$ is randomly sampled from a larger list of 51 types of bugs $\mathbf{B}$ and asks the model to fix issues ranging from state management to code clarity. We generate this list of 51 bugs from papers that study bug distributions in software systems (Just et al., 2014; Widyasari et al., 2020). We intentionally leave the prompt vague to widen the range of acceptable changes. We rollout for a maximum of 115 steps, although this limit is rarely reached.
@@ 3.2 Training
We prioritize training on trajectories that are $\leq$32K tokens in length. To increase sample size as needed, we selectively truncate longer trajectories based on the ratio of trajectory steps within the context limit—we term this “truncation ratio”. In Section 5.2, we explore the effects of truncation in depth and explain why it must be done with caution.
@@ 4.1 Controlled Comparisons
We also note an important methodological consideration regarding context management during evaluation. Some agent frameworks employ optimizations such as retaining only the last few tool calls in context rather than the full trajectory history. While this compression allows models to appear effective at longer context lengths, it introduces a confounding factor: methods using such optimizations may appear to benefit from increased context without the associated computational cost. Furthermore, such optimizations cause key-value cache invalidation during inference, which is prohibitively expensive for practical deployment. For fair comparison, we evaluate all methods using full context retention without such optimizations, ensuring that reported context lengths accurately reflect the actual information available to the model.
@@ 4.3 Repository Specialization
Initial Data Generation: Because every SWE-Bench instance is sourced from a unique commit, the set of instances from each repository will span multiple years. To account for this, we identify the earliest commit and latest commit in SWE-Bench Verified from each repository and generate data from five equally spaced commits in that period. While some functions are repeated across commits, each commit presents the codebase in a different context, which ensures trajectories in both the first and second rollout remain unique. Aggregating across commits, we obtain between 46,000 and 54,000 trajectories for each repository combined across both rollouts. We decide to train on both rollouts to increase sample size since the majority of the generated trajectories exceed 32K tokens. We investigate the effects of mixing rollouts in Section 5.5. Due to compute constraints, we train on 8,000 trajectories per repository rather than the full dataset; however, we release all generated trajectories to enable future research to explore larger-scale specialization.
@@
Table 3: Specialization results at 32K context comparing GLM-4.5-Air (teacher) and fine-tuned Qwen 3-32B (student) on the three largest repositories in SWE-Bench Verified. Fine-tuned models are trained on 8,000 synthetic trajectories from each repository. Results averaged over three seeds. Devstral-Small-2-24B results from Rastogi et al. (2025). See Table 13 in the appendix for 64K evaluation results.
@@
Specialization Scaling Law: To understand how data composition affects specialization efficiency, we fit scaling laws across different mixtures of Django-specific and general coding data (Figure 4). We define the specialization ratio $\alpha$ as the fraction of repository-specific data in the training mixture. At $\alpha=1.0$ (pure Django data), the model matches teacher performance (GLM-4.5-Air at 51.2%) with only 8,000 samples. In contrast, $\alpha=0.0$ (pure general data) is unable to reach equivalent performance even at 16,000 samples. Intermediate mixtures ($\alpha=0.75$, $\alpha=0.25$) show increasing asymptotic performance as the proportion of specialized data increases. This indicates that when training for a target codebase, the ratio of specialized data is the most important factor. A one-way ANOVA revealed a statistically significant effect of specialization ratio at 1,500 samples, $F(3,8)=10.78$, $p=.003$, $\eta^{2}=.80$. Post-hoc comparisons using Tukey’s HSD showed significant differences between $\alpha=1.0$ vs. $\alpha=0.0$ ($p=.009$) and $\alpha=0.75$ vs. $\alpha=0.0$ ($p=.005$). This confirms that repository-specific data yields a statistically significant improvement over general data at equivalent sample sizes.
@@ 5.1 Verification
In Figure 5, we ablate four different verification thresholds for $\mathbf{T_{2}}$ trajectories: $r=0.0$, $0.25$, $0.75$, and $1$. We also plot the performance of models trained on $\mathbf{T_{1}}$ trajectories (which are inherently unverified) at each scale for comparison. We study each verification threshold using only complete trajectories that fit within 32K tokens, with the final datapoint of each curve representing the largest possible dataset for that threshold. Only using trajectories $\leq$32K tokens reduces the total number of trainable trajectories but allows us to avoid confounding factors introduced by truncation.
If verification was essential for performance, we would expect to observe increasing performance as the verification threshold increases. Instead, scaling up to 7,400 samples, we find that all verification thresholds perform similarly. For example, at the maximum scale, training on $\mathbf{T_{2}}$ trajectories that are soft-verified at $r=0.5$ shows no benefit over training on completely unverified $\mathbf{T_{2}}$ trajectories. Furthermore, models trained on $\mathbf{T_{1}}$ trajectories from the first rollout result in similar SWE-Bench Verified performance at each scale, despite representing a completely different distribution of coding tasks.
These results suggest that verification is not a necessity for high quality coding data, a behavior that is similar to what has been observed in other types of reasoning tasks (Chandra et al., 2026). We hypothesize that this is because even incorrect trajectories are can contain important skills, such as how to convert an intention into a relevant code edit, even if the intention does not perfectly address the PR at hand. Indeed, a Kruskal-Wallis H-test revealed no statistically significant difference between verification thresholds, $H(3)=7.19$, $p=.066$, $\varepsilon^{2}=.52$. This indicates that soft verification performs as well as hard verification, and even unverified data achieves comparable results.
@@ 5.2 Truncation
Truncation is a popular but understudied practice across coding agent research to handle trajectories from teacher models whose context windows are larger than the base model’s. Without truncation, base model context limits make it impossible to train on significant amounts of data: $23.26\%$ of SWE-smith and $24.83\%$ of Sera-4.5A-Lite (verified at $r=1$) exceed 32K tokens.
@@
To test this hypothesis, we order $\mathbf{T_{1}}$ trajectories from Sera-4.5A-Lite based on the ratio of trajectory steps that fit in 32K tokens, a property we term “truncation ratio”. We partition the ordered $\mathbf{T_{1}}$ trajectories into subsets of 3,000 samples each. This forces each subsequent partition to contain trajectories with strictly lower truncation ratios than the previous partition. We then train Qwen3-32B on each partition. $\mathbf{T_{1}}$ trajectories work well because they are longer than $\mathbf{T_{2}}$ trajectories on average while exhibiting similar scaling trends. This allows us to study the effect of training on a wide range of truncation ratios with a non-trivial amount of data and expect findings to translate.
In Figure 6 we plot SWE-Bench Verified performance against the average truncation ratio from every partition. Surprisingly, we find that the best data comes from trajectories that have high truncation ratios but are not fully contained in 32K tokens. Subsequent truncation ratios result in gradually decreasing performance. We suspect that this is due to a combination of factors, such as longer trajectories reflecting more difficult tasks and that a model’s final steps are typically focused on the redundant task of submitting its solution instead of problem solving.
@@
Assumptions were met for an independent samples t-test, which showed that ordered truncation ($M=43.00$, $SD=1.93$) significantly outperforms random truncation ($M=37.47$, $SD=0.50$), $t(4)=4.81$, $p=.009$, $d=3.93$. This confirms that the order in which content is truncated matters: preserving earlier turns in the trajectory is more effective than random truncation.